Petty House operates 52 weeks per year, 6 days per week, and uses a continuous review inventory system (i.e. Q system). It purchases kitty litter for $ 10 per bag. The following information is available about these bags. Demand = 102 bags/ week Order cost = $ 49/ order Annual holding cost = 25 percent of cogs Desired cycle service level = 80 percent Lead time = 3 weeks Standard deviation of weekly demand = 15 bags Current on-hand inventory is 320 bags, with no open orders or backorders. a. What is the EOQ? What would be the average time between orders (in weeks)? b. What should R (i.e. reorder point) be? c. An inventory withdrawal of 10 bags was just made. Is it time to reorder? d. The store currently uses a lot size of 600 bags (i. e., Q = 600). What is the annual holding cost of this current policy? Annual ordering cost? Without considering the EOQ, how can you conclude from these two calculations that the current lot size is too large? e. What would be the annual cost saved by shifting from the 600- bag lot size to the EOQ?
Answer #1

Given data,

Days of operations = 6 days per week / 52 Weeks = 312 Days

Demand per week =d = 102 ==> demand per year = 102*52 = 5304 days = D

Order Cost = C = $ 49/order

Price kitty liter = P = $10/bag

Annual Holding Cost = H = 25% of COGS = 0.25* P = 0.25 * 10 = 2.5 $

a) EOQ = SQRT (2* D*C/ H) = SQRT (2*5304*49/2.5) = 455.97= 456 bags

Average Time between orders = EOQ/D = 456/5304 = 0.0859 Year = 4.47 weeks

 

b) Given Standard Deviation = s = 15 = Sigmat

Lead time = 3 weeks

Service level desired = 80%

Reorder point = R = Demand Serviced in order to not lose sales during lead time + Safety stock

Demand Serviced in order to not lose sales during lead time = 102 * 3 = 306

safety stock = z SigmaL = > SigmaL = Sigmat* sqrt(lead time) = 15*sqrt(3) = 25.98

 

for 80% service level , Z = 0.85 ==> safety stock = 306+ (25.8*0.85) = 328

 

 

c) Inventory withdrawal = 10 bag

==> 328 - 10 = 318 bags

Since inventory post withdrawal is less than reorder point, reorder is required

 

d) Annual holding cost with lot size = 600 bags = Q ==> = QH/2 = 600*2.5/2 = $750

Annual ordering cost = Cannual = DC/Q = 5304* 49/600 = $433.16

 

Here annual ordering cost is lesser than holding cost which means current lot size is very large

 

e) Current annual cost with EOQ = (QH/2) + (DC/Q) = 456*2.5/2 + (5304* 49/456) = $ 1139.94

Total savings = 1183.16 – 1139.94 = $43.22

 


answered by: Subrahmanyam golla